OpenStudy (anonymous):

How do I create a quadratic equation for these solutions? x= -5 x= 0

OpenStudy (anonymous):

x^2 + 5x

hero (hero):

x(x+5) = 0

OpenStudy (anonymous):

if you have x=a and x=b as the solutions, then the quadratic equation can be made using (x-a)(x-b)=0

OpenStudy (amistre64):

a quadratic factors to: (x+A) (x+B); where A and B are the opposite values you have

hero (hero):

zero doesn't have an opposite

OpenStudy (amistre64):

..... bummer, what if you turn them over?

OpenStudy (amistre64):

one trick i do, is i give my paper to the person next to me and let them do it ;)

OpenStudy (anonymous):

lol

OpenStudy (mathsolver):

zero doesn't have an opposite so you don't have to add anything to it, the end result would be x(x+5)=0